The specific heat of a metal was found to be 0.38
Answer #1

Let the formula of the metal chloride be MClx

As per Dulong's Petits law,

specific heat*atomic mass of the metal = 25

thus, atomic mass of the metal = 65.79 g/mole

Now, moles of metal in 0.5007 g of it = mass/molar mass = 0.5007/65.79 = 0.0076

mass of Cl in the metal chloride = 1.0586 - 0.5007 = 0.5579

Thus, moles of Cl in the chloride sample = 0.5579/35.5 = 0.0157

Thus, M & Cl are present in the molar ratio of 0.0076:0.0157 = 1:2

Thus, the formula of the compound is MCl2

Now, moles of M in the compound = (1/2)*moles of Cl = (1/2)*0.0157 = 0.00785

Thus, molar mass of M = mass/moles = 0.5007/0.00785 = 63.78 g/mole